OpenStudy (anonymous):

The volume of a rectangular prism is 288 ft3. If its height is 6 feet and its length is 8 feet, what is its width? 6 feet 8 feet 9 feet 12 feet

Directrix (directrix):

Volume of a rectangular solid = l*w*h 288 = 8*6*w 288/ 48 = w 6 feet = w